OpenStudy (anonymous):

a line has a gradient of 3 and passes through the point (-3,3). what is its y intercept?

OpenStudy (anonymous):

start with the point - slope formula \[y-y_1=m(x-x_1)\] put \(m=3,x_1=-3,y_1=3\) and get \[y-3=3(x+3)\] solve for \(y\) to make it look like \(y=mx+b\) and your \(y\)- intercept will be \(b\)

OpenStudy (anonymous):

sorry, i still dont quite get it. so what is the answer? thanks =)

OpenStudy (anonymous):

Well, as satellite73 has already outlined the answer, you just have to re-write the equation as: \[y=3x+12\]. This gives b = 12 which is the y-intercept. To quickly check whether the answer is correct, substitute x = -3 & y=3 in the equation above. It gives an answer 3 = -9 + 12 ==> 3 = 3 Hope it helps..